OpenStudy (anonymous):

re-write the quantity 827,000,000,000,000 picoseconds to show: 4 sig. figs.

OpenStudy (anonymous):

1 sig. figs.=8 x 10^14 2 sig. figs.= 8.3 x 10^14 3 sig. figs. = 8.27 x 10^14 4 sig. figs. = 8.270 x 10^14

OpenStudy (anonymous):

what about 5 sig figs. thats the only one i need

OpenStudy (anonymous):

My apologies, you only asked for 4 sig. figs.. 5 sig. figs 8.2700 x 10^14

OpenStudy (anonymous):

i know my bad but thanks u helped me a lot

OpenStudy (anonymous):

Your Welcome :)